Circular No. 8131

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ef IN NGC 4708 Further to IAUC 8120, D. Weisz and W. Li report the LOTOSS discov-

ery of an apparent supernova on an unfiltered KAIT image taken on May 11.3 UT (mag ∼ 16.3); the new object was confirmed in bright moonlight on May 12.3. SN 2003ef is located at α = 12^{h}49^{m}42^{s}.25, δ = −11^{o}05′29′′.5 (equinox 2000.0), which is 11′′.6 east and 5′′.1 north of the nucleus of NGC 4708. A KAIT image taken on Mar. 22.3 showed nothing at this position (limiting mag ∼ 18.5).

V4743 SAGITTARII M. Orio, Italian National Institute for Astrophysics at Turin, and Uni-

versity of Wisconsin; E. Leibowitz, Tel Aviv University; and P. Rodriguez, European Space Agency, report (on behalf of a team including V. Burwitz, R. Gonzalez-Riestra, Y. Lipkin, J. U. Ness, S. Starrfield, and M. Still) that V4743 Sgr was observed as a target-of-opportunity with the XMM-Newton satellite beginning on Apr. 4.921 UT for ∼ 10 hr: “The EPIC-pn count rate in ‘timing mode’ was 1348.0 ± 0.3 counts/s in the range 0.2–10 keV. The average absorbed flux the range 0.2–2 keV was ∼ 1.5 × 10^{−9} erg cm^{−2} s^{−1}, as high as was observed in the first half of the previous x-ray observation with Chandra (IAUC 8107). The spectrum was extremely soft with absorption features, which we attribute to the central white-dwarf atmosphere. Preliminary results from the lightcurve analysis reveal that the peak in the power spectrum reported on IAUC 8107 is now resolved into two separate peaks, P_{1} = 1309.5 s and Q_{2} = 1375 s. The first overtone of P_{1} is also present in the power spectrum. The period Q_{2} seems to be the first overtone of a fundamental periodicity, P_{2} = 2746 s, that appears significantly in the power spectrum, although with much smaller power. A third period, P_{3} = 1656 s, is probably also present. Despite variability by up to ∼ 20 pecent, the x-ray flux never decreased to the level observed at the end of the recent Chandra observation (IAUC 8107). If an eclipse caused the obscuration observed with Chandra, the orbital period must be longer than 10 hr. We urge optical photometry to detect the orbital period, in order to understand the mechanism that obscured the strongest x-ray source in Mar. 2003.”

CCD V magnitudes measured by J. D. West, Mulvane, KS: 2002 Sept. 24.094 UT, 6.40; 29.073, 7.31; Oct. 21.060, 9.11; 2003 Mar. 21.587, 11.34; 26.574, 11.48; Apr. 4.573, 11.54; 5.570, 11.58; 7.538, 11.49; 9.528, 11.54; 14.523, 11.54; 17.530, 11.69; 21.419, 11.62; 26.553, 11.34; 29.428, 11.68.

2003 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green